# On the parameterized complexity of computing tree-partitions

### v4

@@ 3.3 Time/quality trade-offs
First, we initialize the data structure to compute separators of [BDD+16]. Then we use this data structure to compute separators. We will compute only $k^{O(1)}$ separators per bag, each in time $2^{O(k)}\log n$ which takes $2^{O(k)}\log n$ time per bag. If $alpha=1$, we fully process subinstances and obtain a running time of $2^{O(k)}n\log n$. Otherwise, we stop processing subinstances once they have size $O(\log n)$. If we stop processing subinstances when they reach size $O(\log n)$, we compute only $O(n/\log n)$ bags because the tree-partition is balanced see [BDD+16, Lemma 5.3 and Claim 5.4]. We then compute new tree decompositions for each of the $O(\log n)$ size components in total time $2^{O(k)}n$. For each of them, we apply the algorithm with $\alpha^{\prime}=\alpha-1$, the total running time is then bounded by $2^{O(k)}n\log^{(\alpha^{\prime})}\log n=2^{O(k)}n\log^{(\alpha)}n$.
By fixing $\alpha=\log^{*}n$, we obtain an algorithm running in time $O(2^{O(k)}n\log^{*}n)$.
@@
There is an algorithm running in time $2^{O(k\log k)}n+2^{O(k)}n\log^{*}n$ that computes a tree-partition of width $O(k^{5})$ or reports that the tree-partition-width is more than $k$.
@@
The guarantees on the width follow from the analysis of our scheme (see 3.1 and 3.2).
@@ 4 XALP-completeness of Tree-Partition-Width
XALP is closed by reductions using at most $f(k)\log n$ space and running in FPT time. These two conditions are implied by using at most $f(k)+O(\log n)$ space. We call reductions respecting the latter condition parameterized logspace reductions (or pl-reductions).
@@
Proof 4.2.
To keep things simple, we will use as a black box the fact that reachability in undirected graphs can be decided in logspace [Rei08]. We assume that the vertices have some arbitrary ordering $\sigma$.
@@
We begin by guessing at most $k$ vertices to form an initial bag $B_{0}$, and have an empty parent bag $P_{0}$ We will recursively extend a partial tree-partition in the following manner. Suppose that we have a bag $B$ with parent bag $P$, we must find a child bag for $B$ in each connected component of $G-B$ that does not contain a vertex of $P$. We use the fact that a connected component can be identified by its vertex appearing first in $\sigma$, that the restriction of $\sigma$ to these representatives gives an ordering on $\sigma$, and that we can compute such representatives in logspace. Let us denote by $c$ the current vertex representative of a connected component of $G-B$. $c$ is initially the first vertex in $\sigma$ that is not in $B$ and cannot reach $P$ in $G-B$. We do a co-nondeterministic step so that in one branch of the computation we find a tree-partition for the connected component with representative $c$, and in the other branch we find the representative of the next connected component. The representative $c^{\prime}$ of the next component is the first vertex in $\sigma$ such that it cannot reach a vertex appearing before $c$ (inclusive) in $\sigma$, nor a vertex of $P$ in $G-B$. When found, $c$ is replaced by $c^{\prime}$ and we repeat this computation. If we don’t find such a vertex $c^{\prime}$, then $c$ must have represented the last connected component, so we simply accept.
Let us now describe what happens in the computation branch where we compute a new bag. We can iterate on vertices in the component of $c$, by iterating on vertices of $G-B$ and then skipping if they are not reachable from $c$ in $G-B$. In particular, we can guess a subset $B^{\prime}$ of size at most $k$ of vertices from this component. We then check that the neighbourhood of $B$ in this component is contained in $B^{\prime}$. If it is the case, we can set $P:=B$ and $B:=B^{\prime}$ and recurse. If not, we reject.
@@
During these computations, we store at most $3k+O(1)$ vertices and use logspace subroutines. Furthermore, the described computation tree is of polynomial size.
The following problem is shown to be XALP-complete in [BGJ+22] and is the starting point of our reduction.
Tree-Chained Multicolor Independent Set Input: A tree $T=(V_{T},E_{T})$, an integer $k$, and for each $i\in V_{T}$, a collection of $k$ pairwise disjoint sets of vertices $V_{i,1},\ldots,V_{i,k}$ and a graph $G$ with vertex set $V=\bigcup_{i\in V_{T},j\in[1,k]}V_{i,j}$ Parameter: $k$ Question: Is there a set of vertices $W\subset V$, such that $W$ contains exactly one vertex from each $V_{i,j}$ ($i\in V_{T},j\in[1,k]$), and for each pair $V_{i,j},V_{i^{\prime},j^{\prime}}$ with $i=i^{\prime}$ or $ii^{\prime}\in E_{T}$, $j,j^{\prime}\in[1,k]$, $(i,j)\neq(i^{\prime},j^{\prime})$, the vertex in $W\cap V_{i,j}$ is non-adjacent to the vertex in $W\cap V_{i^{\prime},j^{\prime}}$?
We remark that in the definition above, $V_{i,j}\cap V_{i^{\prime},j^{\prime}}=\emptyset$ whenever $(i,j)\neq(i^{\prime},j^{\prime})$. We further use that we can assume the tree $T$ to be binary without loss of generality (see [BGJ+22] for more details). See Figure 1 for a graphical representation of how the instance is arranged locally.
Figure 1: Local structure of Tree-Chained Multicolor Independent Set. For each $ab\in E_{T}$, the subgraph of $G$ induced by $V_{a}\cup V_{b}$ is a Multicolor Independent Set instance.
We first give a brief sketch of the structure of the hardness proof. We have a trunk gadget to enforce the shape of the tree from the Tree-Chained Multicolor Independent Set. On the trunk are attached clique chains which are longer than the part of trunk between their endpoints, and have some wider parts at some specific positions. The length of the chain gives us some slack which will be used to encode the choice of a vertex for some subset $V_{i,j}$. Based on the edges of $G$, we adjust the width along the trunk so that only one clique chain may place its wider part on each position of the trunk. In other words, part of the trunk is a collection of gadgets representing edges of $G$ that allow for only one incident vertex to be chosen. See Figure 2 for a high level graphical representation of the gadgets.
@@
Proof 4.4.
We reduce from Tree-chained Multicolor Independent Set.
@@
In the Tree-chained Multicolor Independent Set problem, we want to choose one vertex from each set $V_{i,j}$, $i\in V_{T}$, $j\in[1,k]$, such that for each edge $ii^{\prime}\in E_{T}$, the chosen vertices in $V_{i}\cup V_{i^{\prime}}$ form an independent set (which thus will be of size $2k$).
@@
Proof 4.6.
There must be two adjacent bags that contain the vertices of $C_{Z}$ and no other vertices. Similarly, there must be two adjacent bags containing all vertices in $Z\cup\{c_{Z,1},\dots,c_{Z,L}\}$. This forces all vertices in $\{c_{Z,1},\dots,c_{Z,L}\}$ to be in a single bag, and all vertices in $Z$ to be in a single adjacent bag.
@@
The first step in the construction is to build a tree $T^{\prime}=(V_{T^{\prime}},E_{T^{\prime}})$, as follows. Choose an arbitrary node $i$ from $V_{T}$. Add a new neighbour $i^{\prime}$ to $i$, Add a new neighbour $r_{0}$ to $i^{\prime}$. Now subdivide each edge $N$ times. The resulting tree is $T^{\prime}=(V_{T^{\prime}},E_{T^{\prime}})$. We view $T^{\prime}$ as a rooted tree, with root $r_{0}$. We will use the word grandparent to refer to the parent of the parent of a node. The nodes that do not result from the subdivisions are referred to as original nodes (i.e. $V_{T}\cup\{i^{\prime},r_{0}\}$ is the set of original nodes). Nodes $i\in V_{T}$ and their copies in $T^{\prime}$ will be denoted with the same name.
@@
Lemma 4.7.
$H$ has tree-partition-width at most $L$, if and only if the given instance of Tree-chained Multicolor Independent Set has a solution.
Proof 4.8.
Suppose we have a solution of the Tree-chained Multicolor Independent Set. We define $h(i,c)$ for each class $V_{i,c}$, such that the vertex $v_{i,c,h(i,c)}$ is the vertex chosen by the solution in this class. Now, we can construct the tree partition as follows. First, we take the tree $T^{\prime}$, and for each node $i$ in $T^{\prime}$, we take a bag initially containing the vertices in $A_{i}$; we later add more vertices to these bags in the construction.
@@
For a clique chain of $V_{i,c}$, there is a clique mapped to the trunk bag of $i$. Suppose $CC_{i,c,h(i,c)}$ is mapped to $i$. We claim that choosing from each $V_{i,c}$ the vertex $v_{i,c,h(i,c)}$ gives an independent set, and thus, we have a solution of the Tree Chained Multicolor Independent Set problem.
@@
Now, if both endpoints were to be chosen, then the size of the bag of $i_{e_{j}}$ would be larger than $L$: it contains $A_{i_{e_{j}}}$ (which has size $L-6kp(i_{e_{j}})-1$), $kp(i_{e_{j}})$ cliques from clique chains, of which all have size at least 6 and two have size 7; contradiction. So, at most one endpoint is chosen, so choosing vertices $v_{i,c,h(i,c)}$ gives an independent set.
@@
  • A vertex in a trunk clique $A_{i}$ of a node $i$ that is an original node (i.e., also in $T$) has less than $L$ neighbours in $A_{i}$, less than $3L$ neighbours in $A_{i^{\prime}}$ with $i^{\prime}$ incident to $i$ in $T^{\prime}$, less than $kL$ neighbours of cliques $CC_{i,c,1}$ (one clique of size $L-7$ for each class $c\in[1,k]$), less than $4kL$ neighbours of cliques $CC_{i^{\prime},c,2N+r+5}$ (one clique of size $L-7$ for each node of which $i$ is the grandparent in $T$ for each class $c\in[1,k]$), and less than $L$ neighbours in the cluster gadget attached to $A_{i}$.
Finally, we conclude that the transformation can be carried out in $f(k)\log n$ space, thus the result follows.
@@
Proof 4.11.
Membership: We use the fact that the maximum degree of the graph is bounded by $2k$ where $k$ is the domino treewidth. We can discard an instance where this condition on the maximum degree is not satisfied in logspace. We first assume that the given graph is connected.
@@
  • For each neighbour of the bag, whether it was already covered by a bag. This requires $O(k^{2})$ bits (We use some fixed ordering on the vertices to decode which bit corresponds to which vertex in logspace).
The algorithm works as follows. Given the current certificate, if all neighbours have been covered we accept. Otherwise, we guess a new child bag by picking a non-empty subset of $k+1$ vertices among the vertices of the current bag that were contained only in this bag, and the neighbours that were not already covered. We then check that each vertex that is in both the current bag and child bag has all of its neighbours in these two bags. We then guess for each not already covered neighbour of the current bag if it should be covered by the subtree of this child. These vertices are then considered as covered in the current bag certificate. In the child bag certificate, the non covered neighbours are these vertices and the neighbours of the child bag that are not neighbours of the parent bag. We then recurse with both certificates, and accept if both recursions accept.
@@
We can handle disconnected graphs by iterating on component representatives and discarding vertices that are not reachable using the fact that reachability in undirected graphs can be computed in logspace (see membership for Tree-Partition-Width for more details).
@@
First, consider a vertex $w$ in some $C_{v}$. The vertex $y_{w}$ has degree $22M-$ 2, which implies that there are two adjacent bags that each contain $y_{w}$, and $M-1$ neighbours of $y_{w}$. One of these bags contains $w$.
For each $v\in V$, there must be at least one bag that contains all vertices of $C_{v}$, by a well known property of tree decompositions. There can be also at most one such bag, because each vertex $w\in C_{v}$ is in another bag that is filled by $w$, $y_{w}$, and $M-2$ other neighbours of $y_{w}$.
@@
Now, suppose $G$ has tree-partition-width at most $k$, say with tree partition ($(\{Y_{i}|i\in I\},T=(I,F))$. For each $i\in I$, let $X_{i}=\bigcup_{v\in Y_{i}}C_{v}\cup\{z_{e}\leavevmode\nobreak\ |\leavevmode\nobreak\ \exists v\in Y_{i},w\in V:e=vw\}$. For each $v\in V$, $w\in C_{v}$, add two bags, one containing $w$, $y_{w}$, and $M-2$ other neighbours of $y_{w}$, and the other containing $y_{w}$ and the remaining $M-1$ neighbours of $y_{w}$, and make these bags adjacent, with the first adjacent to the bag in $T$ that also contains $w$. One easily verifies that this results in a domino tree decomposition of $H$ with maximum bag size at most $M$, hence $H$ has domino treewidth at most $M-1$.
It is easy to see that $H$ can be constructed from $G$ with $f(k)\log|V|$ memory. So, the hardness of Domino Treewidth follows from the previous lemma.

### v6

@@ 3.3 Time/quality trade-offs
First, we initialize the data structure to compute separators of [7]. Then we use this data structure to compute separators. We will compute only $k^{O(1)}$ separators per bag, each in time $2^{O(k)}\log n$, which takes $2^{O(k)}\log n$ time per bag. If $\alpha=1$, we fully process subinstances and obtain a running time of $2^{O(k)}n\log n$. Otherwise, we stop processing subinstances once they have size $O(\log n)$. If we stop processing subinstances when they reach size $O(\log n)$, we compute only $O(n/\log n)$ bags because the tree-partition is balanced see [7, Lemma 4.3 and Claim 4.4]. We then compute new tree decompositions for each of the $O(\log n)$ size components in total time $2^{O(k)}n$. For each of them, we apply the algorithm with $\alpha^{\prime}=\alpha-1$, the total running time is then bounded by $2^{O(k)}n\log^{(\alpha^{\prime})}\log n=2^{O(k)}n\log^{(\alpha)}n$. ∎
The above algorithm can be turned into an algorithm running in time $2^{O(k)}n$ using the following observation: the number of different configurations to be solved for components of size $O(\log\log(n))$ is small enough to allow us to enumerate all distinct configurations in sublinear time.
Claim 3.9.
We can enumerate all configurations of pairs $(H,S)$ with $H$ a graph on $O(\log\log(n))$ vertices, $S$ a subset of vertices of $H$, compute a tree-partition of $H$ having a superset of $S$ as its root bag for each configuration, and tabulate all results in time $o(n)$ in some table $T$.
Proof.
The number of pairs $(H,S)$ is $2^{O((\log\log(n))^{2})}\cdot 2^{O(\log\log(n))}=2^{O((\log\log(n))^{2})}$. Let $n^{\prime}$ be the number of vertices of $H$. If $k\geq n^{\prime}$, we may use a single bag containing all vertices of $H$. Otherwise, for each such pair, we can compute a tree-partition using the algorithm running in time $2^{O(k)}n^{\prime}\log n^{\prime}$. Using the fact that $k\leq n^{\prime}$, we can bound this running time by $2^{O(n^{\prime})}n^{\prime}\log n^{\prime}=2^{O(n^{\prime})}=2^{O(\log\log(n))}$. We can then store the tree-partition of $H$ which has size $O(n^{\prime})$ in table entry $T[H,S]$.
The total computation time and space is bounded by $2^{O((\log\log(n))^{2})}=o(n)$
Lemma 3.10.
A balanced tree-partition of width $O(\Delta w)$ can be computed in time $2^{O(k)}n$ for a graph of maximum degree $\Delta=k^{O(1)}$ when given a tree decomposition of width $w=O(k)$.
Proof.
We first compute suitable decompositions for all pairs $(H,S)$ of graphs on $O(\log\log(n))$ vertices and subsets of vertices as described in Claim 3.9. We then recursively decompose similarly to Lemma 3.8 with $\alpha=2$ except that now when reaching components of size $O(\log\log(n))$, we simply have to look at a relevant entry of table $T$ to obtain a decomposition for the component.
This is done as follows. We first compute an arbitrary bijection of the vertex set of the component with $[1,n^{\prime}]$ where $n^{\prime}$ is the size of the component. It is then straightforward to obtain an adjacency list describing the component and the list of vertices with neighbors outside the component, both using the new indices. Using this description, we obtain a pair $(H,S)$ allowing us to lookup a decomposition at entry $T[H,S]$. We then obtain a tree-partition with the new indices, and only have to replace all indices by the original indices of vertices in the component. This procedure takes time linear in the size of the component.
The overall running time of the algorithm is $2^{O(k)}n$, since we spend $n+o(n)$ time in total for components of size $O(\log\log(n))$, $2^{O(k)}n\log(n)/\log(n)=2^{O(k)}n$ to reduce the graph to components of size $O(\log(n))$, and $2^{O(k)}(\sum|C_{i}|\log|C_{i}|/\log|C_{i}|)=2^{O(k)}n$ to reduce each component $C_{i}$ of size $O(\log(n))$ to components of size $O(\log\log(n))$. ∎
@@
There is an algorithm running in time $2^{O(k\log k)}n$ that computes a tree-partition of width $O(k^{5})$ or reports that the tree-partition-width is more than $k$.
@@
The guarantees on the width follow from the analysis of our scheme (see 3.1 and 3.2). ∎
@@ 4 XALP-completeness of Tree-Partition-Width
XALP is closed by reductions using at most $f(k)\log n$ space and running in FPT time. These two conditions are implied by using at most $f(k)+O(\log n)$ space. We call reductions respecting the latter condition parameterized log-space reductions (or pl-reductions).
@@
Proof.
To keep things simple, we will use as a black box the fact that reachability in undirected graphs can be decided in log-space [30]. We assume that the vertices have some arbitrary ordering $\sigma$.
@@
We begin by guessing at most $k$ vertices to form an initial bag $B_{0}$, and have an empty parent bag $P_{0}$ We will recursively extend a partial tree-partition in the following manner. Suppose that we have a bag $B$ with parent bag $P$, we must find a child bag for $B$ in each connected component of $G-B$ that does not contain a vertex of $P$. We use the fact that a connected component can be identified by its vertex appearing first in $\sigma$, that the restriction of $\sigma$ to these representatives gives an ordering on $\sigma$, and that we can compute such representatives in log-space. Let us denote by $c$ the current vertex representative of a connected component of $G-B$. $c$ is initially the first vertex in $\sigma$ that is not in $B$ and cannot reach $P$ in $G-B$. We do a co-nondeterministic step so that in one branch of the computation we find a tree-partition for the connected component with representative $c$, and in the other branch we find the representative of the next connected component. The representative $c^{\prime}$ of the next component is the first vertex in $\sigma$ such that it cannot reach a vertex appearing before $c$ (inclusive) in $\sigma$, nor a vertex of $P$ in $G-B$. When found, $c$ is replaced by $c^{\prime}$ and we repeat this computation. If we don’t find such a vertex $c^{\prime}$, then $c$ must have represented the last connected component, so we simply accept.
Let us now describe what happens in the computation branch where we compute a new bag. We can iterate on vertices in the component of $c$, by iterating on vertices of $G-B$ and then skipping if they are not reachable from $c$ in $G-B$. In particular, we can guess a subset $B^{\prime}$ of size at most $k$ of vertices from this component. We then check that the neighborhood of $B$ in this component is contained in $B^{\prime}$. If it is the case, we can set $P:=B$ and $B:=B^{\prime}$ and recurse. If not, we reject.
@@
During these computations, we store at most $3k+O(1)$ vertices and use log-space subroutines. Furthermore, the described computation tree is of polynomial size. ∎
The following problem is shown to be XALP-complete in [9] and is the starting point of our reduction.
Tree-Chained Multicolored Independent Set Input: A tree $T=(V_{T},E_{T})$, an integer $k$, and for each $i\in V_{T}$, a collection of $k$ pairwise disjoint sets of vertices $V_{i,1},\ldots,V_{i,k}$ and a graph $G$ with vertex set $V=\bigcup_{i\in V_{T},j\in[1,k]}V_{i,j}$ Parameter: $k$ Question: Is there a set of vertices $W\subset V$, such that $W$ contains exactly one vertex from each $V_{i,j}$ ($i\in V_{T},j\in[1,k]$), and for each pair $V_{i,j},V_{i^{\prime},j^{\prime}}$ with $i=i^{\prime}$ or $ii^{\prime}\in E_{T}$, $j,j^{\prime}\in[1,k]$, $(i,j)\neq(i^{\prime},j^{\prime})$, the vertex in $W\cap V_{i,j}$ is non-adjacent to the vertex in $W\cap V_{i^{\prime},j^{\prime}}$?
We remark that in the definition above, $V_{i,j}\cap V_{i^{\prime},j^{\prime}}=\emptyset$ whenever $(i,j)\neq(i^{\prime},j^{\prime})$. We further use that we can assume the tree $T$ to be binary without loss of generality (see [9] for more details). See Figure 1 for a graphical representation of how the instance is arranged locally.
Figure 1: Local structure of Tree-Chained Multicolored Independent Set. For each $ab\in E_{T}$, the subgraph of $G$ induced by $V_{a}\cup V_{b}$ is a Multicolored Independent Set instance.
We first give a brief sketch of the structure of the hardness proof. We have a trunk gadget to enforce the shape of the tree from the Tree-Chained Multicolored Independent Set. On the trunk are attached clique chains which are longer than the part of trunk between their endpoints, and have some wider parts at some specific positions. The length of the chain gives us some slack which will be used to encode the choice of a vertex for some subset $V_{i,j}$. Based on the edges of $G$, we adjust the width along the trunk so that only one clique chain may place its wider part on each position of the trunk. In other words, part of the trunk is a collection of gadgets representing edges of $G$ that allow for only one incident vertex to be chosen. See Figure 2 for a high level graphical representation of the gadgets.
@@
Proof.
We reduce from Tree-chained Multicolored Independent Set.
@@
In the Tree-chained Multicolored Independent Set problem, we want to choose one vertex from each set $V_{i,j}$, $i\in V_{T}$, $j\in[1,k]$, such that for each edge $ii^{\prime}\in E_{T}$, the chosen vertices in $V_{i}\cup V_{i^{\prime}}$ form an independent set (which thus will be of size $2k$).
@@
Proof.
There must be two adjacent bags that contain the vertices of $C_{Z}$ and no other vertices. Similarly, there must be two adjacent bags containing all vertices in $Z\cup\{c_{Z,1},\dots,c_{Z,L}\}$. This forces all vertices in $\{c_{Z,1},\dots,c_{Z,L}\}$ to be in a single bag, and all vertices in $Z$ to be in a single adjacent bag. ∎
@@
The first step in the construction is to build a tree $T^{\prime}=(V_{T^{\prime}},E_{T^{\prime}})$, as follows. Choose an arbitrary node $i$ from $V_{T}$. Add a new neighbor $i^{\prime}$ to $i$, Add a new neighbor $r_{0}$ to $i^{\prime}$. Now subdivide each edge $N$ times. The resulting tree is $T^{\prime}=(V_{T^{\prime}},E_{T^{\prime}})$. We view $T^{\prime}$ as a rooted tree, with root $r_{0}$. We will use the word grandparent to refer to the parent of the parent of a node. The nodes that do not result from the subdivisions are referred to as original nodes (i.e. $V_{T}\cup\{i^{\prime},r_{0}\}$ is the set of original nodes). Nodes $i\in V_{T}$ and their copies in $T^{\prime}$ will be denoted with the same name.
@@
Claim 4.4.
$H$ has tree-partition-width at most $L$, if and only if the given instance of Tree-chained Multicolored Independent Set has a solution.
Proof.
Suppose we have a solution of the Tree-chained Multicolored Independent Set. We define $h(i,c)$ for each class $V_{i,c}$, such that the vertex $v_{i,c,h(i,c)}$ is the vertex chosen by the solution in this class. Now, we can construct the tree partition as follows. First, we take the tree $T^{\prime}$, and for each node $i$ in $T^{\prime}$, we take a bag initially containing the vertices in $A_{i}$; we later add more vertices to these bags in the construction.
@@
For a clique chain of $V_{i,c}$, there is a clique mapped to the trunk bag of $i$. Suppose $CC_{i,c,h(i,c)}$ is mapped to $i$. We claim that choosing from each $V_{i,c}$ the vertex $v_{i,c,h(i,c)}$ gives an independent set, and thus, we have a solution of the Tree Chained Multicolored Independent Set problem.
@@
Now, if both endpoints were to be chosen, then the size of the bag of $i_{e_{j}}$ would be larger than $L$: it contains $A_{i_{e_{j}}}$ (which has size $L-6kp(i_{e_{j}})-1$), $kp(i_{e_{j}})$ cliques from clique chains, of which all have size at least 6 and two have size 7; contradiction. So, at most one endpoint is chosen, so choosing vertices $v_{i,c,h(i,c)}$ gives an independent set. ∎
@@
  • A vertex in a trunk clique $A_{i}$ of a node $i$ that is an original node (i.e., also in $T$) has less than $L$ neighbors in $A_{i}$, less than $3L$ neighbors in $A_{i^{\prime}}$ with $i^{\prime}$ incident to $i$ in $T^{\prime}$, less than $kL$ neighbors of cliques $CC_{i,c,1}$ (one clique of size $L-7$ for each class $c\in[1,k]$), less than $4kL$ neighbors of cliques $CC_{i^{\prime},c,2N+r+5}$ (one clique of size $L-7$ for each node of which $i$ is the grandparent in $T$ for each class $c\in[1,k]$), and less than $L$ neighbors in the cluster gadget attached to $A_{i}$.
Finally, we conclude that the transformation can be carried out in $f(k)\log n$ space, thus the result follows. ∎
@@
Proof.
We first prove membership in XALP.
We use the fact that the maximum degree of the graph is bounded by $2k$ where $k$ is the domino treewidth. We can discard an instance where this condition on the maximum degree is not satisfied in log-space. We first assume that the given graph is connected.
@@
  • For each neighbor of the bag, whether it was already covered by a bag. This requires $O(k^{2})$ bits (We use some fixed ordering on the vertices to decode which bit corresponds to which vertex in log-space).
The algorithm works as follows. Given the current certificate, if all neighbors have been covered we accept. Otherwise, we guess a new child bag by picking a non-empty subset of $k+1$ vertices among the vertices of the current bag that were contained only in this bag, and the neighbors that were not already covered. We then check that each vertex that is in both the current bag and child bag has all of its neighbors in these two bags. We then guess for each not already covered neighbor of the current bag if it should be covered by the subtree of this child. These vertices are then considered as covered in the current bag certificate. In the child bag certificate, the non covered neighbors are these vertices and the neighbors of the child bag that are not neighbors of the parent bag. We then recurse with both certificates, and accept if both recursions accept.
@@
We can handle disconnected graphs by iterating on component representatives and discarding vertices that are not reachable using the fact that reachability in undirected graphs can be computed in log-space (see membership for Tree-Partition-Width for more details).
@@
First, consider a vertex $w$ in some $C_{v}$. The vertex $y_{w}$ has degree $22M-$ 2, which implies that there are two adjacent bags that each contain $y_{w}$, and $M-1$ neighbors of $y_{w}$. One of these bags contains $w$.
For each $v\in V$, there must be at least one bag that contains all vertices of $C_{v}$, by a well known property of tree decompositions. There can be also at most one such bag, because each vertex $w\in C_{v}$ is in another bag that is filled by $w$, $y_{w}$, and $M-2$ other neighbors of $y_{w}$.
@@
Now, suppose $G$ has tree-partition-width at most $k$, say with tree partition ($(\{Y_{i}|i\in I\},T=(I,F))$. For each $i\in I$, let $X_{i}=\bigcup_{v\in Y_{i}}C_{v}\cup\{z_{e}~{}|~{}\exists v\in Y_{i},w\in V:e=vw\}$. For each $v\in V$, $w\in C_{v}$, add two bags, one containing $w$, $y_{w}$, and $M-2$ other neighbors of $y_{w}$, and the other containing $y_{w}$ and the remaining $M-1$ neighbors of $y_{w}$, and make these bags adjacent, with the first adjacent to the bag in $T$ that also contains $w$. One easily verifies that this results in a domino tree decomposition of $H$ with maximum bag size at most $M$, hence $H$ has domino treewidth at most $M-1$. ∎
It is easy to see that $H$ can be constructed from $G$ with $f(k)\log|V|$ memory. So, the hardness of Domino Treewidth follows from the previous lemma. ∎